### v1

@@
Multi-District School Choice: Playing on Several Fieldsthanks: This work adapts and extends the undergraduate thesis of Yin (2022). Gonczarowski gratefully acknowledges research support by the Harvard FAS Inequality in America Initiative and by the Harvard FAS Dean’s Competitive Fund for Promising Scholarship. Zhang was supported by an NSF Graduate Research Fellowship.
@@
   Michael Yin Department of Applied Mathematics, Harvard University. Email: michaelyin2018@gmail.com.    Shirley Zhang Department of Computer Science, Harvard University. Email: szhang2@g.harvard.edu.
We extend the seminal model of Pathak and Sönmez (2008) to a setting with multiple school districts, each running its own separate centralized match, and focus on the case of two districts. In our setting, in addition to each student being either sincere or sophisticated, she is also either constrained—able to apply only to schools within her own district of residence—or unconstrained—able to choose any single district within which to apply. We show that several key results from Pathak and Sönmez (2008) qualitatively flip: A sophisticated student may prefer for a sincere student to become sophisticated, and a sophisticated student may prefer for her own district to use Deferred Acceptance over the Boston Mechanism, irrespective of the mechanism used by the other district. We furthermore investigate the preferences of students over the constraint levels of other students. Many of these phenomena appear abundantly in large random markets.
@@ 1 Introduction
The Boston Mechanism (henceforth BM; also sometimes referred to as the “Immediate Acceptance” mechanism) is a widely used school-choice mechanism, especially in school-choice systems that were never (re)designed by economists or computer scientists. This mechanism first maximizes the number of applicants who get their first-choice school (breaking ties based on the priorities that students have at the different schools, i.e., based on the schools’ “preferences”); then, subject to that, maximizes the number of applicants who get their second-choice school; then, subject to that, maximizes the number of applicants who get their third-choice school; and so forth. Despite being a very natural mechanism, BM suffers from various unattractive qualities, such as not being strategyproof and resulting in unstable matchings. Due to these and other shortcomings, there has been a push since the turn of the millenium (Abdulkadiroğlu et al., 2005b, a; Pathak and Sönmez, 2008) to replace BM with the better-behaved Deferred Acceptance mechanism (Gale and Shapley, 1962, henceforth DA) in school-choice systems.
@@
One of the most compelling arguments given in favor of replacing BM with DA is the equity argument that originates in the seminal paper of Pathak and Sönmez (2008), which considers a setting with some students being sincere (i.e., uninformed and unstrategic, always reporting their true preferences) and some being sophisticated (i.e., informed and strategic, together playing a Nash equilibrium). That paper proves that sophisticated students weakly prefer BM over DA (which could be seen as a baseline that treats sincere and sophisticated students equally, due to its strategyproofness). This leads Pathak and Sönmez (2008) to view BM as weakly (and many times strictly) conferring an advantage to sophisticated students over sincere ones. Pathak and Sönmez (2008) furthermore prove that when BM is used, sophisticated students weakly prefer for sincere students to remain sincere, giving a plausible explanation as to why informed parent groups might not be likely to share their know-how with parents outside their groups or social circles.
@@
A school district running a centralized matching mechanism is not an isolated capsule. Many districts, each running an independent centralized match, might exist next to each other, and some students might be able to effectively choose which district’s match to participate in. For example, a 2005 report for the Berkeley Unified School District in California (Fried, 2005) estimated that between 7.8% and 12% of the district’s high schoolers were “attending [the district] unofficially”, and actually lived out-of-district. Choosing one’s school district can thus be done without official permission (as in the case above) at personal risk, [^1] See, for instance, Martin (2011) for more context on this illegal phenomenon, known as “boundary hopping” or “residency fraud,” which at times has led to a prison sentence for parents. or legally by moving to that district, an option many times available only to populations with greater financial resources. [^2] Comparing the prices of houses located near school district boundaries, Black (1999) estimates that parents are willing to pay 2.1% more to enroll their child in a district with a 5% higher mean test score, and Bayer et al. (2007) similarly estimate a 1.8% higher willingness to pay for homes in a district with an average test score that is higher by one standard deviation.
@@
Neighbouring school districts are often independent of each other and might use different mechanisms. Due to the ability of some students to choose their school districts, one district switching its mechanism has the potential to change the multi-district equilibrium, changing students’ strategies not only in terms of how they rank schools within a school district but also in terms of their choice of school district. In this paper, we examine the two predictions of Pathak and Sönmez (2008) that we describe above in a multi-district setting in which district choice (by the students who, for instance, possess the resources to officially relocate or are willing to risk punishment) is endogenized as part of the equilibrium. That is, in our setting, in addition to each student being either sincere or sophisticated, she is also either constrained—able to apply only to schools within her own district of residence—or unconstrained—able to strategically choose any single district within which to apply. [^3] In either case, if she is sophisticated, she can strategically order her submitted preference list over the schools in the district to which she applies. We prove that even when considering only two school districts, both of the predictions of Pathak and Sönmez (2008) that we describe above flip. That is, a sophisticated student may strictly prefer for her district to use DA over BM, irrespective of whether she is constrained or unconstrained and of the mechanism used by the other district. Furthermore, a sophisticated student may strictly prefer for some sincere student to become sophisticated. The latter phenomenon also appears abundantly in large random markets, that is, as the size of the market grows, a constant fraction of sophisticated students strictly prefer that at least some sincere students become sophisticated. Finally, we complete our investigation by asking whether some students might prefer for others to change their constraint type, e.g., whether an unconstrained student might prefer for another student to become unconstrained, or whether a constrained student might prefer for another student who resides in a different district to become constrained. We prove a strong “anything goes” result showing that every possible such combination is abundant in a large random market.
Our results are not without limitations. For one, consider the phenomena of sophisticated students strictly preferring DA over BM. While we show this phenomenon to be possible, it might be rare in random markets, [^4] We do show that its frequency at the very least does not diminish as the market grows. which could still lend credence to an argument in favor of DA over BM. Importantly, though, this argument becomes a quantitative issue of relative frequency rather than a qualitative issue of existence. Pathak and Sönmez (2008), in their study of the one-district setting, speculate that school-choice discussion organizations like the West Zone Parents Group may exist to coordinate behavior among sophisticated students without informing sincere ones. While our results indicate that in the multi-district setting it is feasible for such organizations to have incentives to inform some sincere students and increase their sophistication, these groups do not necessarily have a means of targeting specifically the sincere students who improve some sophisticated students’ outcomes, again rendering the story less clear cut.
Our paper, therefore, is not intended to advocate for the use of BM. Rather, first and foremost, it serves to introduce a formal model of multi-district school choice and highlight that taking into account the broader landscape beyond only a single district may qualitatively change the analysis, including the arguments for or against the use of various mechanisms. Specifically, our results provide a proof-of-concept that highlights that district choice, which manifests as the real-life behavior of sending one’s child to school in a desirable district by either using a false address or paying a premium to move, plays an important role in shaping the playing field for both sophisticated and sincere students. Our results should be interpreted as “anything goes” results, highlighting that when designing the specifics of a mechanism or market (even when choosing, for instance, between different ways to break ties in students’ priorities at schools if these priorities are not strict), there are no one-size-fits-all solutions. Instead, one must weigh the specifics of the market in question, perhaps even more broadly defined than usually considered.
@@ 1.1 Related Work
To our knowledge, ours is the first theoretical analysis of multi-district school choice. Closest to our work are previous papers that analyze different types of schools (such as charter, magnet, and private schools) coexisting with public schools. For instance, considering schools that are not district-run and can therefore choose to use their own admissions systems, Ekmekci and Yenmez (2019) analyze school incentives for participating in a unified enrollment system in a single district. Other papers assume there is no option for a unified enrollment system, and instead analyze a “slightly decentralized” mechanism that can be used to rematch students with the vacant seats that arise from schools of several types accepting the same student (Manjunath and Turhan, 2016; Turhan, 2019; Afacan et al., 2022). Akbarpour et al. (2022) study how exogenously varying the value of outside options affects behavior and preference over mechanisms. Although these threads of research bear some resemblance to the multi-district school choice problem that we study, there are a number of key differences. Ekmekci and Yenmez (2019) fix the students in the district and focus on the choice by schools of whether to participate in a unified enrollment system; we instead fix the schools in each of multiple districts and allow some students to choose in which district to apply. Meanwhile, Manjunath and Turhan (2016), Turhan (2019), and Afacan et al. (2022) allow every student to report rankings for each school type and choose between matches they receive; Akbarpour et al. (2022) similarly do not require students to forego their outside option in order to participate in a match. By contrast, we distinguish between students who can and cannot utilize district choice, and even those who can are only able to rank schools in their single chosen district (and therefore only receive one match, and cannot keep a guaranteed outside option). Our paper centers on how students endogenously choose their district of enrollment and thus affect the landscape of a school choice problem, a modeling decision that distinguishes this paper from the above prior work.
@@ 2.1.2 Mechanisms
  • Round 1: Each student who submitted a non-empty ROL applies to the school she ranked 1st on her ROL. For each school, if there are at least as many seats available as applicants, the school (permanently) accepts every applicant. Otherwise, each school allocates seats to applicants based on the school’s priority ordering up to its capacity, and rejects the remaining students for whom no seats remain.
  • Round $k>1$: Consider only students who have not yet been accepted to a school (i.e., the students who have been rejected by the schools $1$st through $(k\!-\!1)$th on their ROLs). Each student who submitted an ROL of at least length $k$ applies to the school she ranked $k$th. For each school, applicants are accepted or rejected in the same way as in Round 1, where the seats available are those that were not already filled in previous rounds. If a school has no seats available at the beginning of the round, it rejects all new applicants.
@@
  • Round 1: Each student who submitted a non-empty ROL applies to the school ranked 1st on her ROL. For each school, if there are at least as many seats available as applicants, the school tentatively accepts every applicant. Otherwise, each school tentatively allocates seats to applicants based on the school’s priority ordering up to its capacity, and (permanently) rejects the remaining students for whom no seats remain.
  • Round $k>1$: Consider only students who are not currently tentatively accepted at a school (i.e., the students who were rejected by a school in round $k\!-\!1$). Each of these students applies to the school highest on her ROL that has not already rejected her. For each school, new applicants are considered alongside tentatively accepted students. All of these students are compared based on the school’s priority ordering and are tentatively accepted or permanently rejected in the same way as in Round 1, where all seats at the school are initially considered available. [^5] Students who were tentatively accepted by the school in round $k\!-\!1$ are not conferred any advantage, and may still be permanently rejected by that school in round $k$.
@@
We say that a mechanism is strategyproof if truthful reporting is a dominant strategy for every student. BM is not strategyproof: a student may benefit from reporting an ROL that differs from her true preference ordering. DA is strategyproof (Dubins and Freedman, 1981; Roth, 1982): regardless of other students’ reported ROLs, it is every student’s dominant strategy to report her true preference ordering as her ROL.
@@ 2.2 Multi-District School Choice
In this paper, we extend the traditional school choice problem (henceforth, the single-district school choice problem) by considering multiple districts. Specifically, in a multi-district school choice problem, there is a set of districts numbered $\{1,...,\ell\}$. Each student $i$ resides in some district $d(i)$ and each school $s$ is located in some district $d(s)$. A student’s preference ordering may be over schools in multiple districts (including districts in which the student does not reside), and a school’s priority ordering is over all students across all districts. A student’s ROL may only contain schools from a single district, however; we call this the district in which she enrolls. Intuitively, this models the real-world setting wherein a student can only enroll in a single school district in a given year.
@@
Each district uses its own school choice mechanism to determine school assignments of the students who enroll in the district to the schools that are located in the district. Different districts may use the same mechanism or different mechanisms; regardless, the school assignments for each district are combined to form the school assignments for the multi-district school choice problem as a whole.
2.3 Student Sophistication Types and Constraints Types
@@
As in Pathak and Sönmez (2008), a student is either sincere or sophisticated; this is known as her sophistication type. Once a sincere student $i$ determines she will enroll in some district $j$, she submits her preference ordering limited to schools in $j$ (i.e., $P_{i}$ with any schools not in $j$ removed) as her ROL. In other words, a sincere student reports her true preferences over schools in the district she enrolls in. On the other hand, a sophisticated student can strategize by submitting any ROL over schools in the district she enrolls in. In addition to having a sophistication type, in the multi-district setting, a student is also either constrained or unconstrained; we refer to this as her constraint type. A constrained student $i$ can only enroll in the district in which she resides, $d(i)$, while an unconstrained student can enroll in any (single) district.
@@
Combining these two attributes, we have four categories of students: sincere-constrained, sincere-unconstrained, sophisticated-constrained, and sophisticated-unconstrained. The behavior of each of these is largely intuitive, with a sincere-constrained student reporting her true preferences over schools in her district of residence; a sophisticated-constrained student strategically choosing an ROL over schools in her district of residence; and a sophisticated-unconstrained student strategically choosing both a district to enroll in and an ROL to submit over schools in that district.
One might consider two different definitions for sincere-unconstrained students. A sincere-unconstrained student could enroll in the district of her first-choice school and then report her true preferences over schools in that district; such a student does not strategize at all. Alternatively, a sincere-unconstrained student could also know that she will report her true preferences over schools in whichever district she enrolls in, but strategically choose which district to enroll in. By the first definition, a sincere-unconstrained student who can never get assigned to her first-choice school will still enroll in the district of that school; in contrast, using the second definition, a sincere-unconstrained student would instead enroll in another district if she would get a preferable school assignment by reporting her true preferences over schools in that district. We believe that both of these definitions are have merit in different contexts, so we ensure that our results hold regardless of which definition is used. Where relevant, we demonstrate in our examples and theorems that the outcome is the same whether the first or second definition of sincere-unconstrained students is used.

### v2

@@
Multi-District School Choice: Playing on Several Fieldsthanks: This work adapts and extends the undergraduate thesis of Yin (2022). We thank Scott Kominers and Assaf Romm for insightful comments and discussions. Gonczarowski gratefully acknowledges research support by the National Science Foundation (NSF-BSF grant No. 2343922), Harvard FAS Inequality in America Initiative, and Harvard FAS Dean’s Competitive Fund for Promising Scholarship. Zhang was supported by an NSF Graduate Research Fellowship.
@@
   Michael Yin Paris School of Economics. Email: michaelyin2018@gmail.com.    Shirley Zhang Department of Computer Science, Harvard University. Email: szhang2@g.harvard.edu.
We extend the seminal model of Pathak and Sönmez (2008) to a setting with multiple school districts, each running its own separate centralized match, and focus on the case of two districts. In our setting, in addition to each student being either sincere or sophisticated, she is also either constrained—able to apply only to schools within her own district of residence—or unconstrained—able to choose any single district within which to apply. We show that several key results from Pathak and Sönmez (2008) no longer hold: A sophisticated student may prefer for a sincere student to become sophisticated, and a sophisticated student may prefer for her own district to use Deferred Acceptance over the Boston Mechanism, irrespective of the mechanism used by the other district. We furthermore investigate the preferences of students over the constraint levels of other students. Many of these phenomena appear abundantly in large random markets.
@@ 1 Introduction
The Boston Mechanism (henceforth, BM; also sometimes referred to as the “Immediate Acceptance” mechanism) is a widely used school-choice mechanism, especially in school-choice systems that were not (re)designed by economists or computer scientists. This mechanism first maximizes the number of applicants who get their first-choice school (breaking ties based on the priorities that students have at the different schools, i.e., based on the schools’ “preferences”); then, subject to that, maximizes the number of applicants who get their second-choice school; then, subject to that, maximizes the number of applicants who get their third-choice school; and so forth. Despite being a very natural mechanism, BM suffers from various unattractive qualities, such as not being strategyproof and resulting in unstable matchings. Due to these and other shortcomings, there has been a push since the turn of the millennium (e.g., Abdulkadiroğlu et al., 2005a, b; Pathak and Sönmez, 2008) to replace BM with the better-behaved Deferred Acceptance mechanism (Gale and Shapley, 1962; henceforth, DA) in school-choice systems.
@@
One of the most compelling arguments given in favor of replacing BM with DA is the equity argument that originates in the seminal paper of Pathak and Sönmez (2008), which considers a setting with some students being sincere (i.e., uninformed and unstrategic, always reporting their true preferences) and some being sophisticated (i.e., informed and strategic, together playing a Nash equilibrium). That paper proves that sophisticated students weakly prefer BM over DA (the latter mechanism could be seen as a baseline that treats sincere and sophisticated students equally, due to its strategyproofness). This leads Pathak and Sönmez (2008) to view BM as weakly (and many times strictly) conferring an advantage to sophisticated students over sincere ones. Pathak and Sönmez (2008) furthermore prove that when BM is used, sophisticated students weakly prefer for sincere students to remain sincere, giving a plausible explanation as to why informed parent groups might not be likely to share their know-how with parents outside their groups or social circles.
@@
A school district running a centralized matching mechanism is not an isolated capsule. Many districts, each running an independent centralized match, might exist next to each other, and some students might be able to choose in which district’s match to participate. For example, a 2005 report for the Berkeley Unified School District in California (Fried, 2005) estimated that between 7.8% and 12% of the district’s high schoolers were “attending [the district] unofficially,” and actually lived out-of-district. Choosing one’s school district can thus be done without official permission (as in the case above) at personal risk, [^1] See, for instance, Martin (2011) for more context on this illegal phenomenon, known as “boundary hopping” or “residency fraud,” which has at times led to prison sentences for parents. or legally by moving to that district, an option many times available only to populations with greater financial resources. [^2] Comparing the prices of houses located near school district boundaries, Black (1999) estimates that parents are willing to pay 2.1% more to enroll their child in a district with a 5% higher mean test score, and Bayer et al. (2007) similarly estimate a 1.8% higher willingness to pay for homes in a district with an average test score that is higher by one standard deviation.
@@
Neighboring school districts are often independent of one another and might use different mechanisms to match students to schools. Due to the ability of some students to choose their school districts, one district switching its mechanism has the potential to change the multi-district equilibrium, changing students’ strategies not only in terms of how they rank schools within a district but also in terms of their choice of school district. In this paper, we examine the two predictions of Pathak and Sönmez (2008) that we describe above in a multi-district setting in which district choice (by the students who, for instance, possess the resources to officially relocate or are willing to risk punishment) is endogenized as part of the equilibrium. That is, in our setting, in addition to each student being either sincere or sophisticated, she is also either constrained—able to apply only to schools within her own district of residence—or unconstrained—able to strategically choose any single district within which to apply. [^3] In either case, if a student is sophisticated, she strategically orders her submitted preference list over the schools in the district to which she applies.
We prove that even when considering only two school districts, both of the predictions of Pathak and Sönmez (2008) that we describe above no longer hold. Specifically, a sophisticated student may strictly prefer for her district to use DA over BM, irrespective of whether she is constrained or unconstrained and of the mechanism used by the other district. Furthermore, a sophisticated student may strictly prefer for some sincere student to become sophisticated. The latter phenomenon also appears abundantly in large random markets; that is, for sufficiently large random markets, a constant fraction of sophisticated students strictly prefer that at least some sincere students become sophisticated. We round out our investigation by asking, for completeness, whether some students might prefer for others to change their constraint types, e.g., whether an unconstrained student might prefer for another student to become unconstrained, or whether a constrained student might prefer for another student who resides in a different district to become constrained. We prove a strong “anything goes” result showing that every possible such combination appears abundantly in large random markets.
Our results are not without limitations. For one, consider the phenomenon of sophisticated students strictly preferring DA over BM. While we show this phenomenon to be possible, it is our only result that might be rare in random markets, [^4] We do show that its frequency at least does not diminish as the market grows. which could still lend credence to an argument in favor of DA over BM. Importantly, though, this argument becomes a quantitative issue of relative frequency rather than a qualitative issue of existence. More broadly, our paper is not intended to advocate for the use of BM. Rather, first and foremost, it serves to introduce a formal model of multi-district school choice and provide a proof-of-concept highlighting that taking into account the broader landscape beyond only a single district may qualitatively change the analysis, including the validity of certain arguments for or against the use of various mechanisms. When designing the specifics of a mechanism or market (be it when choosing the overall mechanism as discussed in this paper, or possibly even when considering far more minute implementation details), one always weighs the specifics of the market in question. Our results highlight that in some cases, this market should be even more broadly defined than is customary.
The remainder of this paper is structured as follows. After reviewing related work, in Section 2 we present the multi-district school choice model. In Section 3 we prove our first main result, regarding preference over sophistication types. In Section 4 we prove our second main result, regarding preference over mechanisms. In Section 5 we prove our results regarding preference over constraint types. In Section 6, we dive deeper into the mechanics that enable our first main result—that sophisticated students might strictly prefer for some or all sincere students to become sophisticated. We uncover that there are two distinct mechanisms that can drive this result, identify the precise features of our model that enable each of these mechanisms, show their robustness, and derive necessary as well as sufficient conditions over the market structure for each of these mechanisms to manifest. We conclude with a discussion in Section 7.
@@ 1.1 Related Work
To our knowledge, ours is the first theoretical analysis of multi-district school choice where some students can choose their district. Hafalir et al. (2022) also discuss a notion of interdistrict school choice, but in their model all students can freely rank schools regardless of district, and the focus is on a policy goal of diversity across districts. Grigoryan (2023) considers multiple neighborhoods with a school in each one, but families can again freely rank schools regardless of neighborhood, and the emphasis is on aggregate welfare and welfare for low-income families under different matching mechanisms. In contrast to these papers, because we utilize the lens of constrained and unconstrained students, our multi-district school choice problems are distinct from large single-district problems, allowing us to re-examine the equity arguments of Pathak and Sönmez (2008) in this setting.
Within a single district, closest to our work are previous papers that analyze different types of schools (such as charter, magnet, and private schools) coexisting with public schools. For instance, considering schools that are not district-run and can therefore choose to use their own admissions systems, Ekmekci and Yenmez (2019) analyze school incentives for participating in a unified enrollment system in a single district. Doğan and Yenmez (2019), meanwhile, compare the efficiency of unified and divided enrollment systems in a district that has selective and nonselective schools (where students can be admitted to a school of each type under the divided enrollment system). Other papers assume that there is no option for a unified enrollment system, and instead analyze a “slightly decentralized” mechanism that can be used to rematch students with the vacant seats that arise from schools of several types accepting the same student (Manjunath and Turhan, 2016; Turhan, 2019; Afacan et al., 2022). Akbarpour et al. (2022) study how exogenously varying the value of outside options affects behavior and preference over mechanisms. Although these threads of research bear some resemblance to the multi-district school choice problem that we study, there are a number of key differences. Ekmekci and Yenmez (2019) fix the students in the district and focus on the choice by schools of whether to participate in a unified enrollment system; we instead fix the schools in each of multiple districts and allow some students to choose in which district to apply. Meanwhile, Manjunath and Turhan (2016), Turhan (2019), and Afacan et al. (2022) allow every student to report rankings for each school type and choose between matches they receive; the same is true for Doğan and Yenmez (2019) under divided enrollment. Akbarpour et al. (2022) similarly do not require students to forego their outside option in order to participate in a match. By contrast, we distinguish between students who can and cannot utilize district choice, and even those who can are only able to rank schools in their single chosen district (and therefore only receive one match, and cannot keep a guaranteed outside option). Our paper centers on how students endogenously choose their district of enrollment and thus affect the landscape of a school choice problem, a modeling decision that distinguishes this paper from the above prior work.
@@ 2.1.2 Mechanisms
  • Round 1: Each student who submitted a non-empty ROL applies to the school she ranked 1st on her ROL. For each school, if there are at least as many seats available as applicants, the school (permanently) accepts every applicant. Otherwise, each school (permanently) allocates seats to applicants based on the school’s priority ordering up to its capacity, and rejects the remaining students for whom no seats remain.
  • Round $k>1$: Consider only students who have not yet been accepted to a school (i.e., the students who have been rejected by the schools 11st through (k1)(k\!-\!1)th on their ROLs). Each student who submitted an ROL of at least length $k$ applies to the school she ranked $k$th. For each school, applicants are accepted or rejected in the same way as in Round 1, where the seats available are those that were not already filled in previous rounds. If a school has no seats available at the beginning of the round, it rejects all new applicants.
@@
  • Round 1: Each student who submitted a non-empty ROL applies to the school she ranked 1st on her ROL. For each school, if there are at least as many seats available as applicants, the school tentatively accepts every applicant. Otherwise, each school tentatively allocates seats to applicants based on the school’s priority ordering up to its capacity, and (permanently) rejects the remaining students for whom no seats remain.
  • Round $k>1$: Consider only students who are not currently tentatively accepted at a school (i.e., the students who were rejected by a school in round k1k\!-\!1). Each of these students applies to the school highest on her ROL that has not already rejected her. For each school, new applicants are considered alongside tentatively accepted students. All of these students are compared based on the school’s priority ordering and are tentatively accepted or permanently rejected in the same way as in Round 1, where all seats at the school are initially considered available. [^5] Students who were tentatively accepted by the school in round k1k\!-\!1 are not conferred any advantage, and may still be permanently rejected by that school in round $k$.
@@
We say that a mechanism is strategyproof if truthful reporting is a dominant strategy for every student. BM is not strategyproof: a student may benefit from reporting an ROL that differs from her true preference ordering. DA is strategyproof (Dubins and Freedman, 1981; Roth, 1982): regardless of other students’ reported ROLs, it is a dominant strategy for every student to report her true preference ordering as her ROL.
@@ 2.2 Multi-District School Choice
In this paper, we extend the traditional school choice problem (henceforth, the single-district school choice problem) by considering multiple districts. Specifically, in a multi-district school choice problem, there is a set of $\ell$ districts. Each student $i$ resides in some district $d(i)$ and each school $s$ is located in some district $d(s)$. A student’s preference ordering may be over schools in multiple districts (including districts in which the student does not reside), and a school’s priority ordering is over all students across all districts. A student’s ROL may only contain schools from a single district, however; we call this the district in which she enrolls. Intuitively, this models the real-world setting wherein a student can only enroll in a single school district in a given year.
@@
Each district uses its own school choice mechanism to determine assignments of the students who enroll in the district to the schools that are located in the district. Different districts may use the same mechanism or different mechanisms; regardless, the school assignments for each district are combined to form the school assignments for the multi-district school choice problem as a whole.
2.3 Student Sophistication Types and Constraint Types
@@
As in Pathak and Sönmez (2008), a student is either sincere or sophisticated; this is known as her sophistication type. Once a sincere student $i$ determines that she will enroll in some district $j$, she submits her preference ordering restricted to schools in $j$ (i.e., $P_{i}$ with any schools not in $j$ removed) as her ROL. In other words, a sincere student reports her true preferences over schools in the district she enrolls in. On the other hand, a sophisticated student strategizes when submitting her ROL over schools in the district she enrolls in. In addition to having a sophistication type, in the multi-district setting a student is also either constrained or unconstrained; we refer to this as her constraint type. A constrained student $i$ can only enroll in the district in which she resides, $d(i)$; while an unconstrained student can enroll in any (single) district.
@@
Combining these two attributes, we have four categories of students: sincere-constrained, sincere-unconstrained, sophisticated-constrained, and sophisticated-unconstrained. The behavior of each of these is largely intuitive, with a sincere-constrained student reporting her true preferences over schools in her district of residence; a sophisticated-constrained student strategically choosing an ROL over schools in her district of residence; and a sophisticated-unconstrained student strategically choosing both a district to enroll in and an ROL to submit over schools in that district. As in Pathak and Sönmez (2008), sophisticated students who enroll in a DA district always use their dominant strategy within that district, i.e., truthfully rank the schools in that district. If an unconstrained student is indifferent between districts to enroll in (since each student has a strict preference ordering over schools she finds acceptable, such indifference can only occur if this student remains unassigned regardless of where she enrolls), then we assume that this student enrolls in her district of residence.
For the majority of this paper, we consider a sincere-unconstrained student to be one who reports her true preferences over schools in whichever district she enrolls in, and strategically chooses in which district to enroll accordingly. This is not the only reasonable definition of sincere-unconstrained students, and in Section 6.2 we ensure that our results hold for a large class of other reasonable definitions as well; see Section 6 for a more in-depth discussion of sincere-unconstrained students.